I am not sure how to figure out this equation, can you help explain step by step?
The endpoints of the diameter of a circle are (-7,1) and the center point of x^+y^-2x-10y+25=0. Find the standard equation of this circle and graph. Find and label the x-intercepts and the y-intercepts.
I appreciate any help you can give.

Answer by venugopalramana(3286)

I am not sure how to figure out this equation, can you help explain step by step?
The endpoints of the diameter of a circle are (-7,1) and the center point of x^+y^-2x-10y+25=0. Find the standard equation of this circle and graph. Find and label the x-intercepts and the y-intercepts.
I appreciate any help you can give.
I THINK YOU MEAN.....
LET S= x^2+y^2-2x-10y+25=0.
PLEASE TYPE PROPERLY.
CENTRE OF IS OBTAINED BY
(X-1)^2+(Y-5)^2-1-25+25=0
CENTRE IS (1,5)...WITH SAME CENTRE BUT DIFFERENT RADIUS EQN IS GIVEN BY
(X-1)^2+(Y-5)^2=R^2....IT IS PASSSING THROUGH(-7,1)...SO
(-7-1)^2+(1-5)^2=R^2
64+16=R^2=80
HENCE EQN.OF REQD.CIRCLE IS
(X-1)^2+(Y-5)^2=80
X INTERCEPTS ARE GOT BY PUTTING Y=0
(X-1)^2=80-25=55...X-1=+-SQRT(55)
X=1+-SQRT(55)
Y INTERCEPTS ARE OBTAINED BY PUTTING X=0
(Y-5)^2=80-1=79
Y=5+-SQRT(79)